Fermi National Laboratory


Accelerator Update for January 27, 2006 - January 30, 2006

Friday January 27

The day shift began with Tevatron (Tev) in access for B4 vacuum investigation, the Recycler Ring (RR) stash at 21.9 E10, Antiproton Source (Pbar) stack at 81.4 E10, NuMI and MCenter (MIPP) off due to Main Injector (MI) power supply problem, and with MiniBooNE taking beam.

MI beam returns at 10:08 AM.

Beam returns to NuMI, MIPP and Meson Test (MT) at 10:23 AM.

Operations put the Recycler into shot setup at 10:52 AM to stash antiprotons.

Pbar returns to stacking at 11:59 AM, beam also returns to MCenter.

At 1:23 PM, the Tevatron B4 access is finished, turn on is underway.

Operations put the Recycler into shot setup at 4:59 PM to stash antiprotons.

Beam returns to MCenter at 6:02 PM, Pbar remains off to investigate a high temperature problem at the APO water cage.

Pbar returns to stacking at 6:22 PM.

Tevatron shot setup begins at 8:36 PM.

Operations established store #4612 at 11:26 PM with an initial luminosity of 53.02 E30. The store was lower than expected due to problems with the TEL.

Saturday January 28

The midnight shift began with Operations monitoring store #4612, the Recycler Ring (RR) stash at 2.9 E10, Pbar stack at 59.1 E10, and with MiniBooNE, NuMI, and MCenter taking beam.

Pbar stacking went down at 12:42 AM due to problems with a septa (D:ISEPV).

Operations put the Recycler into shot setup at 1:14 AM to stash antiprotons.

Beam returns to MCenter at 2:23 AM, Pbar remains off due to the septa problem.

Pbar stacking resumes at 3:27 AM.

Operations put the Recycler into shot setup at 8:47 AM to stash antiprotons.

Pbar returns to stacking at 9:34 AM, beam also returns to MCenter.

MT turns on at 1:00 PM.

Operations put the Recycler into shot setup at 5:50 PM to stash antiprotons.

Beam returns to MCenter at 6:42 PM, Pbar remains off for access, to investigate problems with the Debuncher low level amplifiers.

Pbar returns to stacking at 7:34 PM.

MT turns off at 10:04 PM.

Store 4612 was terminated at 10:06 PM.

Sunday January 29

The midnight shift began with Tev studies, the Recycler Ring (RR) stashing, with Pbar stacking, and with MiniBooNE, NuMI and MIPP taking beam.

Tevatron shot setup begins at 2:58 AM.

Operations established store #4614 at 4:23 AM with an initial luminosity of 69.1 E30.

The store luminosity was once again lower than expected due to problems with the TEL.

Operations put the Recycler into shot setup at 4:32 AM to stash antiprotons.

Pbar returns to stacking at 5:40 AM, beam also returns to MCenter.

At 8:13 AM the Booster gradient magnet power supply (B:GMPS) tripped off on a ground fault indication, experts are investigating, no proton beam anywhere while this is off.

Booster is turning on at 3:00 PM, the ground fault was caused by a split sump discharge hose spraying on to the magnet.

Beam returns at 3:23 PM.

Monday January 30

The midnight shift began with Operations monitoring store #4614, the Recycler Ring (RR) stashing, with Pbar stacking, and with MiniBooNE, NuMI and MIPP taking beam.

Pbar stacking was halted at 7:30 AM due to beam instabilities.

MIPP turns off at 7:51 AM to repair a water leak on F:MC7AN1.

The Plan for Monday

Put in a low luminosity Tev store, stack, stash and continue to deliver beam to NuMI and MiniBooNE. Turn on MIPP once the MC7AN1 leak is repaired.

 
Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.



last modified 01/30/2006    email Fermilab